# ReLLa: Retrieval-enhanced Large Language Models for Lifelong Sequential Behavior Comprehension in Recommendation

### v3

@@ Abstract.
With large language models (LLMs) achieving remarkable breakthroughs in natural language processing (NLP) domains, LLM-enhanced recommender systems have received much attention and have been actively explored currently. In this paper, we focus on adapting and empowering a pure large language model for zero-shot and few-shot recommendation tasks. First and foremost, we identify and formulate the lifelong sequential behavior incomprehension problem for LLMs in recommendation domains, i.e., LLMs fail to extract useful information from a textual context of long user behavior sequence, even if the length of context is far from reaching the context limitation of LLMs. To address such an issue and improve the recommendation performance of LLMs, we propose a novel framework, namely Retrieval-enhanced Large Language models (ReLLa) for recommendation tasks in both zero-shot and few-shot settings. For zero-shot recommendation, we perform semantic user behavior retrieval (SUBR) to improve the data quality of testing samples, which greatly reduces the difficulty for LLMs to extract the essential knowledge from user behavior sequences. As for few-shot recommendation, we further design retrieval-enhanced instruction tuning (ReiT) by adopting SUBR as a data augmentation technique for training samples. Specifically, we develop a mixed training dataset consisting of both the original data samples and their retrieval-enhanced counterparts. We conduct extensive experiments on three real-world public datasets to demonstrate the superiority of ReLLa compared with existing baseline models, as well as its capability for lifelong sequential behavior comprehension. To be highlighted, with only less than 10% training samples, few-shot ReLLa can outperform traditional CTR models that are trained on the entire training set (e.g., DCNv2, DIN, SIM). The code is available [^1] https://github.com/LaVieEnRose365/ReLLa.
@@ 5.2. Language Models for Recommendation
As suggested in previous work (Lin et al., 2023a), the adaption of language models to the field of recommender systems can be generally categorized according to the roles they serve in the recommendation pipeline, i.e., feature engineering (Liu et al., 2023a; Borisov et al., 2022; Li et al., 2023c; Mysore et al., 2023; Carranza et al., 2023; Christakopoulou et al., 2023), feature encoder (Muhamed et al., 2021b; Hou et al., 2022; Yu et al., 2021; Wang et al., 2022b; Hou et al., 2023a; Zhang et al., 2022; Fu et al., 2023b; Yuan et al., 2023; Qiu et al., 2021; Li et al., 2023b; He et al., 2022; Muhamed et al., 2021b), scoring/ranking function (Liu et al., 2022; Kang et al., 2023; Zhang et al., 2021a; Li et al., 2023e; Bao et al., 2023; Li et al., 2023d; Zhang and Wang, 2023; Mao et al., 2023; Hua et al., 2023a; Geng et al., 2023; Hua et al., 2023b; Zhang et al., 2023a; Hou et al., 2023b; Chen, 2023; Petrov and Macdonald, 2023; Wang and Lim, 2023).
@@ Acknowledgements.
The Shanghai Jiao Tong University team is partially supported by National Key R&D Program of China (2022ZD0114804), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and National Natural Science Foundation of China (62177033, 62322603).

### v6

@@ Abstract.
With large language models (LLMs) achieving remarkable breakthroughs in natural language processing (NLP) domains, LLM-enhanced recommender systems have received much attention and have been actively explored currently. In this paper, we focus on adapting and empowering a pure large language model for zero-shot and few-shot recommendation tasks. First and foremost, we identify and formulate the lifelong sequential behavior incomprehension problem for LLMs in recommendation domains, i.e., LLMs fail to extract useful information from a textual context of long user behavior sequence, even if the length of context is far from reaching the context limitation of LLMs. To address such an issue and improve the recommendation performance of LLMs, we propose a novel framework, namely Retrieval-enhanced Large Language models (ReLLa) for recommendation tasks in both zero-shot and few-shot settings. For zero-shot recommendation, we perform semantic user behavior retrieval (SUBR) to improve the data quality of testing samples, which greatly reduces the difficulty for LLMs to extract the essential knowledge from user behavior sequences. As for few-shot recommendation, we further design retrieval-enhanced instruction tuning (ReiT) by adopting SUBR as a data augmentation technique for training samples. Specifically, we develop a mixed training dataset consisting of both the original data samples and their retrieval-enhanced counterparts. We conduct extensive experiments on three real-world public datasets to demonstrate the superiority of ReLLa compared with existing baseline models, as well as its capability for lifelong sequential behavior comprehension. To be highlighted, with only less than 10% training samples, few-shot ReLLa can outperform traditional CTR models that are trained on the entire training set (e.g., DCNv2, DIN, SIM). The code is available [^1] PyTorch version: https://github.com/LaVieEnRose365/ReLLa [^2] MindSpore version: https://github.com/mindspore-lab/models/tree/master/research/huawei-noah/ReLLa.
@@ 5.2. Language Models for Recommendation
As suggested in previous work (Lin et al., 2023a), the adaption of language models to the field of recommender systems can be generally categorized according to the roles they serve in the recommendation pipeline, i.e., feature engineering (Liu et al., 2023a; Borisov et al., 2022; Li et al., 2023c; Mysore et al., 2023; Carranza et al., 2023; Christakopoulou et al., 2023), feature encoder (Muhamed et al., 2021b; Hou et al., 2022; Yu et al., 2021; Wang et al., 2022b; Hou et al., 2023a; Zhang et al., 2022; Fu et al., 2023b; Yuan et al., 2023; Qiu et al., 2021; Li et al., 2023b), scoring/ranking function (Liu et al., 2022; Kang et al., 2023; Zhang et al., 2021a; Li et al., 2023e; Bao et al., 2023; Li et al., 2023d; Zhang and Wang, 2023; Mao et al., 2023; Hua et al., 2023a; Geng et al., 2023; Hua et al., 2023b; Zhang et al., 2023a; Hou et al., 2023b; Chen, 2023; Petrov and Macdonald, 2023; Wang and Lim, 2023).
@@ Acknowledgements.
The Shanghai Jiao Tong University team is partially supported by National Key R&D Program of China (2022ZD0114804), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and National Natural Science Foundation of China (62177033, 62322603). The work is sponsored by Huawei Innovation Research Program. We thank MindSpore (min, 2020) for the partial support of this work, which is a new deep learning computing framework.